How do you determine the revenue?

Blossoms Agency sells an insurance policy offered by Capital Insurance Company for a commission of $86 on January 2, 2017. Blossom will receive an additional commission of $8 each year for as long as the policyholder does not cancel the policy. After selling the policy, Blossom does not have any remaining performance obligations. Based on Blossoms significant experience with these types of policies, it estimates that policyholders on average renew the policy for 4.5 years after the first year before terminating their insurance policy. It has no evidence to suggest that previous policyholder behavior will change. Determine the transaction price of the arrangement for Blossom, assuming 90 policies are sold. Transaction price $ 1098D ] Determine the revenue that Blossom will recognize in 2017. (Round answer to O decimal places, e.g. 5,125) Revenue
Answer #1
a) Calculation of transaction price
no of policies = 90
Commission on policies = 86
Additional commission = 8
year of renewal =4.5
transaction price = 90*86 + 90*8*4.5 =
transaction price = 7740+3240
therefore transaction price = 10980
b) Revenue that blossom will recognize in 2017 = transaction price/ 4.5
Revenue recognize in 2017 = 10980/4.5
Revenue recognize in 2017 = $ 2440